OpenStudy (anonymous):

a typical person has an average heart beat rate of 69.0 beats per minute. How many beats does she have in 3.0 years?

OpenStudy (anonymous):

First find the total number of minutes in three years. Using dimensional analysis you get 1576800 minutes in three years. (60 mins * 24 hrs * 365 days * 3 years) Then just multiply that number to figure out how many beats an average person has in three years. 108799200 beats in three years.

OpenStudy (kkutie7):

\[\frac{69 beats}{min}*\frac{60 min}{1hr}*\frac{24 hr}{1day}*\frac{3(360)day}{3(1)year}\] \[=\frac{35769600beats}{3years}\]